Need to find the dimension of a rectangle with a perimeter of 640yds. All I have is one side is w and the other is 3w-12. 3w has me confused. Thanks.

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

2w+2(3w-12)=640
2w+6w-24=640
8w=640+24
8w=684
w=684/8
w=85.5
L=(3w-12)=3*85.5-12
L=244.5

Answer by mananth(16946)

Perimeter = 2*(L+W)
640= 2*(W+3W-12))
640=2(4W-12)
640=8W-24
8W=664
/8
W=83 yards one side
other side = 3w-12
3*83-12
249-12
237 yards
219+77=326